OpenStudy (anonymous):

a coffee maker costs $49.95. How much will it cost after a 20% discount?

OpenStudy (anonymous):

PLZ GIVE ANSWER WITH FORMULA AND EXPLANATION THANKS:)

OpenStudy (anonymous):

A 20% discount will mean that the coffee maker will now only cost 80% of the original price. So if we were to multiply the original cost by 1, we would get the same price i.e. 100% of the original cost. Similarly if we multiply the original cost by 0.8, we will get 80% of the original price; \[49.95 \times 0.8 = 39.96\]

OpenStudy (anonymous):

that wasnt what i was looking for but thanks for trying:) @mafffs

OpenStudy (anonymous):

What were you looking for?

OpenStudy (anonymous):

something like this: percent change= to-from ------- from

OpenStudy (anonymous):

\[percentage change = \frac{original percentage - discounted percentage}{original percentage}\] \[discounted price = \frac{original percentage - discounted percentage}{original percentage} \times price\]

OpenStudy (anonymous):

$9.99 is the answer

OpenStudy (anonymous):

$9.99 is the amount taken off the original price to give us the discounted price. The discounted price is what your question was asking for.

OpenStudy (anonymous):

Yeah, sorry I remember now, ahaahah subract

OpenStudy (anonymous):

I'm glad :)